OpenStudy (anonymous):

Find the rectangular coordinates for the point whose polar coordinates are given. (7√2, -π/4)

OpenStudy (amistre64):

x = rcos(t); y = rsin(t) r = 7sqrt(2) ; t = -pi/4

OpenStudy (anonymous):

Thank you! You are the only one who I can actually understand.

OpenStudy (amistre64):

its my southern drawl :)

OpenStudy (anonymous):

I bet. People in the north are very rude :p